# Reparameterized LLM Training via Orthogonal Equivalence Transformation

### v3

@@ 3.1 General Framework
Figure 3: Singular values of a weight matrix of size 512$\times$1376, randomly generated by different initialization schemes.
@@
3.2 Efficient Approximations to Orthogonality
@@
Figure 4: Examples of the primitive orthogonal transformation matrix $\bm{G}_{i}$ in different orthogonalizations (two examples for each method). Note that, blue blocks represent $1$, light purple blocks denote 00 and deep purple blocks are the actual orthogonal parameterization to be learned.
@@ 3.2.1 Stochastic Primitive Optimization
Fully stochastic SPO. The basic idea of fully stochastic SPO is to randomly sample a small submatrix and enforce its orthogonality, allowing it to be easily extended to a full orthogonal matrix by embedding it within an identity matrix–a process similar to Givens or Householder transformations. To represent a large orthogonal matrix $\bm{R}\in\mathbb{R}^{m\times m}$, we start by defining $c$ index sets $\bm{S}^{j}=\{s^{j}_{1},\cdots,s^{j}_{b}\}\subseteq\{1,\cdots,m\}$ ($j\in[1,c]$), where each set has cardinality $|\bm{S}^{j}|=b$, a hyperparameter controlling the number of effective parameters of a primitive orthogonal matrix. $\bm{S}^{j},\forall j$ are randomly sampled from the full indices $\{1,\cdots,m\}$. Let $\tilde{\bm{G}}_{j}\in\mathbb{R}^{b\times b}$ be a small orthogonal matrix, and $\bm{D}(\bm{S}^{j})=\{\bm{e}({s^{j}_{1}}),\cdots,\bm{e}({s^{j}_{b}})\}\in\mathbb{R}^{m\times b}$ be a selection matrix, where $\bm{e}(k)$ is the standard basis vector with a $1$ in the $k$-th position and 00 elsewhere. The factorization is given by
@@ 3.2.2 Cayley-Neumann Parameterization
$\small\bm{R}=(\bm{I}+\bm{Q})(\bm{I}-\bm{Q})^{-1}=(\bm{I}+\bm{Q})\cdot\big{(}\sum_{i=0}^{\infty}\bm{Q}^{i}\big{)}\approx(\bm{I}+\bm{Q})\cdot\big{(}\bm{I}+\sum_{i=1}^{k}\bm{Q}^{i}\big{)},$ (6)
@@ 3.2.3 Overall Training Algorithm
Step 2: Orthogonal matrix initialization. For fully stochastic SPO, we randomly sample an index set $\bm{S}$, and parameterize $\tilde{\bm{G}}_{R}\in\mathbb{R}^{b\times b}$ and $\tilde{\bm{G}}_{P}\in\mathbb{R}^{b\times b}$ using CNP (Equation (6)). Both matrices are initialized as identity, so $\bm{R}$ and $\bm{P}$ also start as identity matrices. For block-stochastic SPO, we sample a random permutation matrix $\bm{\Psi}_{R},\bm{\Psi}_{P}$, and parameterize $\{\tilde{\bm{G}}_{R}^{1},\cdots,\tilde{\bm{G}}_{R}^{\lceil\frac{m}{b}\rceil}\}$ and $\{\tilde{\bm{G}}_{P}^{1},\cdots,\tilde{\bm{G}}_{P}^{\lceil\frac{m}{b}\rceil}\}$ using CNP. Then we initialize them as the identity, so $\bm{R}$ and $\bm{P}$ again starts as identity matrices.
Step 3: Efficient orthogonal parameterization. For fully stochastic SPO, we have $\bm{R}=\bm{I}_{m}+\bm{D}(\bm{S})(\tilde{\bm{G}}_{R}-\bm{I}_{b})\bm{D}(\bm{S})^{\top}$ and $\bm{P}=\bm{I}_{m}+\bm{D}(\bm{S})(\tilde{\bm{G}}_{P}-\bm{I}_{b})\bm{D}(\bm{S})^{\top}$. For block-stochastic SPO, we have $\bm{R}=\bm{\Psi}_{R}^{\top}\text{Diag}(\tilde{\bm{G}}^{1}_{R},\cdots,\tilde{\bm{G}}^{\lceil\frac{m}{b}\rceil}_{R})\bm{\Psi}_{R}$ and $\bm{P}=\bm{\Psi}_{P}^{\top}\text{Diag}(\tilde{\bm{G}}^{1}_{P},\cdots,\tilde{\bm{G}}^{\lceil\frac{m}{b}\rceil}_{P})\bm{\Psi}_{P}$.
@@ 4 Discussions and Intriguing Insights
AdamW $mn$ $3mn$
GaLore [81] $mn$ $mn+mr+2nr$
POET (FS) $b(b-1)$ $mn+3b(b-1)$
@@
Performance under a constant parameter budget. Since POET optimizes two orthogonal matrices $\bm{R},\bm{P}$ simultaneously, a natural question arises: which matrix should receive more parameter budget under a fixed total constraint? To investigate this, we conduct a controlled experiment where different ratios of trainable parameters are allocated to $\bm{R}$ and $\bm{P}$ under a fixed total budget. All other settings (e.g., architecture, data) remain unchanged, with full details provided in the Appendix. We use validation perplexity as the evaluation metric. The total parameter budget matches that of fully stochastic POET with $b=\frac{1}{h}m$ for $\bm{R}$ and $b=\frac{1}{h}n$ for $\bm{P}$, where $h=8$, $4$, and $3$ correspond to small, medium, and large budgets, respectively. We explore seven allocation settings: $\bm{R}:\bm{P}=1:0$ (i.e., orthogonal training [48, 65, 51]), $0.9:0.1$, $0.75:0.25$, $0.5:0.5$ (i.e., standard POET), $0.25:0.75$, $0.1:0.9$, and $0:1$. Results in Figure 5 show that POET with a balanced allocation between $\bm{R}$ and $\bm{P}$ yields the best performance.
@@
[…]
@@
The proof of this lemma can be found in Appendix D. The result extends to Block-stochastic SPO as this is strictly more expressive than fully stochastic SPO. The key idea of the proof is similar to the factorization of orthogonal matrices into a product of Givens rotations. Indeed, by multiplying $\bm{R}^{\top}$ with properly chosen primitive matrices $\bm{G}_{i}$ we can create zeros below the diagonal for one column after another. Note that each $\bm{G}_{i}$ has $b(b-1)/2$ parameters while $\bm{R}$ has $m(m-1)/2$ parameters, which implies that generally at least $\Omega((m/b)^{2})$ primitives are necessary. In Appendix D we also provide a heuristic that with high probability for $c=O(\ln(m)(m/b)^{2})$ every orthogonal matrix can be written as a product of $c$ orthogonal primitives $\bm{G}_{i}$.
@@
Dynamics of singular spectrum. Inspired by [46], we conduct a spectral analysis by comparing the singular spectrum among AdamW, Muon and POET. Following [1, 68], we compute SVD entropy of the trained Llama-60M model at different iteration. Specifically, the SVD entropy, defined as $H(\bm{\sigma})=\frac{1}{\log n}\sum_{i}\frac{\sigma_{i}^{2}}{\sum_{j}\sigma_{j}^{2}}\log\frac{\sigma_{i}^{2}}{\sum_{j}\sigma_{j}^{2}}$ measures the diversity of singular values; higher entropy indicates a more uniform and diverse spectrum. [46] attributes the favorable performance of Muon over AdamW to the more diverse spectrum of weight matrices updates. As shown in Figure 6, POET consistently maintains high spectral diversity throughout training, owing to its orthogonal equivalence transformation. Therefore, POET can better explore diverse optimization directions. Consistent with the observation in [46], we also find that Muon yields more diverse spectral updates than AdamW.
POET minimizes hyperspherical energy in expectation. While POET effectively generalizes energy-preserving training to spectrum-preserving training, it still ensures low hyperspherical energy when initialized with a zero-mean isotropic Gaussian distribution. This is actually a significant property, as POET retains the generalization benefits of minimal hyperspherical energy [47, 49, 48]. This property comes from the invariance of zero-mean isotropic Gaussian to orthogonal transformation. [48] shows that for a weight matrix $\bm{W}$ that is initialized by zero-mean isotropic Gaussian, its neurons, after being normalized, are uniformly distributed on the unit hypersphere. This property provably leads to a small hyperspherical energy. Since zero-mean isotropic Gaussian initialization is invariant to orthogonal transformation, $\text{OET}(\bm{W};\bm{R},\bm{P})=\bm{R}\bm{W}\bm{P}$ does not change the distribution of $\bm{W}$, i.e., $\bm{R}\bm{W}\bm{P}=_{d}\bm{W}$. We give a detailed derivation in Appendix B. Therefore, the transformed neurons of the equivalent weight matrix $\bm{R}\bm{W}\bm{P}$ are also uniformly distributed over the unit hypersphere after being normalized. This validates that POET can provably preserve a small energy.
@@
The property that POET preserves singular spectrum while retaining minimal hyperspherical energy in expectation well justifies why POET yields stable training and good generalization performance. Such a property only holds true when the weight initialization follows zero-mean isotropic Gaussian, which is perfectly satisfied by current weight initialization. It may also justify why zero-mean isotropic Gaussian initialization (e.g., standard Gaussian, normalized Gaussian, Xavier [17]) works better than the uniform spectrum initialization in our experiments. To further validate that POET maintains a small hyperspherical energy, we plot the hyperspherical energy of the Llama-60M model at different iterations in Figure 7. We compare the sum of the hyperspherical energy of all the layers trained by AdamW, Muon or POET. The results again verify POET’s energy-minimizing property. Slightly different from previous orthogonal training [48], POET can not preserve exactly the same hyperspherical energy throughout the training, but it can provably minimize hyperspherical energy in an expectation sense. Interestingly, we find that Muon can also minimize hyperspherical energy more effectively than AdamW, which partially justifies why Muon may lead to better performance.
@@ 5.1 LLM Pretraining using LLaMA Transformers
POETBS,$b$=64 29.52 (2.39M) 24.52 (5.52M) 20.29 (14.90M) 18.28 (29.22M)
POETBS,$b$=128 26.90 (4.81M) 21.86 (11.12M) 18.05 (30.04M) 16.24 (58.91M)
POETBS,$b$=256 25.29 (9.66M) 19.88 (22.33M) 16.27 (60.32M) 14.56 (118.26M)
POETFS,$b$=1/8 34.06 (0.53M) 29.67 (1.78M) 24.61 (6.34M) 18.46 (25.39M)
POETFS,$b$=1/4 28.69 (2.13M) 23.55 (7.13M) 19.42 (25.44M) 17.60 (101.66M)
POETFS,$b$=1/2 25.37 (8.54M) 19.94 (28.56M) 15.95 (101.86M) 13.70 (406.88M)
@@
We perform the pretraining experiments on the Llama transformers of varying sizes (60M, 130M, 350M, 1.3B) for POET. We use the C4 dataset [66], a cleaned web crawl corpus from Common Crawl, widely used for LLM pretraining [81, 56, 29]. For POET-BS, $b$ is the block size of the block-diagonal orthogonal matrix. For POET-FS, $b_{\text{in}}$=$bm$ for $\bm{R}$ and $b_{\text{out}}$=$bn$ for $\bm{P}$. We compare POET against GaLore [81], a low-rank pretraining method, and AdamW, the standard pretraining optimizer. We generally follow the settings in [81]. To better simulate the practical pretraining setting, we significantly increase the number of training tokens for all methods.
@@
Table 2 shows that both POET-FS ($b$=1/2) and POET-BS ($b$=256) consistently outperform both GaLore and AdamW with significantly fewer parameters. For LLaMA-1B, POET-FS ($b$=1/2) yields the best overall performance, achieving a validation perplexity of $13.70$, much better than AdamW ($14.73$) and GaLore ($18.33$). Block-stochastic POET with $b$=256 achieves the second-best performance ($14.56$), which still surpasses AdamW with only one-tenth of AdamW’s trainable parameters. Similar patterns can be observed for models of smaller sizes. Moreover, we compare the training dynamics between AdamW and POET in Figure 8. The training dynamics of POET is quite different from AdamW. After an initial rapid drop in perplexity, POET improves more slowly than AdamW. As seen in Phase II (Figure 2), this slower but stable progress can lead to better performance in later stages. We attribute this intriguing phenomenon to the unique reparameterization of POET and How we efficiently approximate orthogonality. The exact mechanism behind this phenomenon remains an open question, and understanding it could offer valuable insights into large-scale model training.
@@
Figure 9: Validation perplexity dynamics of POET (FS, $b$=1/2) and AdamW on Llama-60M. POET outperforms the AdamW trained with almost twice the number of seen tokens.
@@
To highlight POET’s non-trivial performance improvement, we increase the training steps (i.e., effectively tokens seen) for AdamW, and find that POET-FS ($b$=1/2) still outperforms AdamW even even if AdamW is trained with almost triple the number of tokens. Results are given in Figure 9. In this experiment, the AdamW learning rate was carefully tuned for the full training run, and no training tokens were repeated. Thus, the improvement is non-trivial and cannot be attributed to merely increasing training steps. Interestingly, we also observe from Table 2 that POET’s performance appears strongly correlated with the parameter budget and larger budgets consistently yield better results across model scales. This is particularly important for model scaling law [35]. Another notable observation is that POET significantly outperforms LoRA [26] given a similar parameter budget. For instance, with approximately 30M trainable parameters, POET attains a validation perplexity of $18.05$, significantly better than LoRA’s $25.19$. We further observe that the block-stochastic variant is more parameter-efficient than the fully stochastic one. On the 130M model, it achieves a validation perplexity of $19.88$ with nearly 6M fewer trainable parameters, compared to $19.94$ for the fully stochastic variant. We hypothesize that this is due to better coverage of weight parameters. Specifically, the block-stochastic variant ensures all corresponding weights are updated at each step, unlike the more uneven updates in the fully stochastic variant. Experimental details and results on weight update coverage are provided in Appendix H.

### v4

@@ 3.1 General Framework
[…]
@@
3.2 Efficient Approximation to Orthogonality
@@
Figure 4: Examples of the primitive orthogonal transformation matrix $\bm{G}_{i}$ in different orthogonalizations (two examples for each method). Note that, blue blocks represent 11, light purple blocks denote 0 and deep purple blocks are the actual orthogonal parameterization to be learned.
@@ 3.2.1 Stochastic Primitive Optimization
Fully stochastic SPO. The basic idea of fully stochastic SPO is to randomly sample a small submatrix and enforce its orthogonality, allowing it to be easily extended to a full orthogonal matrix by embedding it within an identity matrix–a process similar to Givens or Householder transformations. To represent a large orthogonal matrix $\bm{R}\in\mathbb{R}^{m\times m}$, we start by defining $c$ index sets $\bm{S}^{j}=\{s^{j}_{1},\cdots,s^{j}_{b}\}\subseteq\{1,\cdots,m\}$ ($j\in[1,c]$), where each set has cardinality $|\bm{S}^{j}|=b$, a hyperparameter controlling the number of effective parameters of a primitive orthogonal matrix. $\bm{S}^{j},\forall j$ are randomly sampled from the full indices $\{1,\cdots,m\}$. Let $\tilde{\bm{G}}_{j}\in\mathbb{R}^{b\times b}$ be a small orthogonal matrix, and $\bm{D}(\bm{S}^{j})=\{\bm{e}({s^{j}_{1}}),\cdots,\bm{e}({s^{j}_{b}})\}\in\mathbb{R}^{m\times b}$ be a selection matrix, where $\bm{e}(k)$ is the standard basis vector with a $1$ in the $k$-th position and 0 elsewhere. The factorization is given by
@@ 3.2.2 Cayley-Neumann Parameterization
$\small\bm{R}=(\bm{I}+\bm{Q})(\bm{I}-\bm{Q})^{-1}=(\bm{I}+\bm{Q})\cdot\big(\sum_{i=0}^{\infty}\bm{Q}^{i}\big)\approx(\bm{I}+\bm{Q})\cdot\big(\sum_{i=0}^{k}\bm{Q}^{i}\big)=(\bm{I}+\bm{Q})\cdot\big(\bm{I}+\sum_{i=1}^{k}\bm{Q}^{i}\big),$ (6)
@@ 3.2.3 Overall Training Algorithm
Step 2: Orthogonal matrix initialization. For fully stochastic SPO, we randomly sample an index set $\bm{S}$, and parameterize $\tilde{\bm{G}}_{R}\in\mathbb{R}^{b\times b}$ and $\tilde{\bm{G}}_{P}\in\mathbb{R}^{b\times b}$ using CNP (Equation (6)). Both matrices are initialized as identity, so $\bm{R}$ and $\bm{P}$ also start as identity matrices. For block-stochastic SPO, we sample a random permutation matrix $\bm{\Psi}_{R},\bm{\Psi}_{P}$, and parameterize $\{\tilde{\bm{G}}_{R}^{1},\cdots,\tilde{\bm{G}}_{R}^{\lceil\frac{m}{b}\rceil}\}$ and $\{\tilde{\bm{G}}_{P}^{1},\cdots,\tilde{\bm{G}}_{P}^{\lceil\frac{n}{b}\rceil}\}$ using CNP. Then we initialize them as the identity, so $\bm{R}$ and $\bm{P}$ again starts as identity matrices.
Step 3: Efficient orthogonal parameterization. For fully stochastic SPO, we have $\bm{R}=\bm{I}_{m}+\bm{D}(\bm{S})(\tilde{\bm{G}}_{R}-\bm{I}_{b})\bm{D}(\bm{S})^{\top}$ and $\bm{P}=\bm{I}_{n}+\bm{D}(\bm{S})(\tilde{\bm{G}}_{P}-\bm{I}_{b})\bm{D}(\bm{S})^{\top}$. For block-stochastic SPO, we have $\bm{R}=\bm{\Psi}_{R}^{\top}\text{Diag}(\tilde{\bm{G}}^{1}_{R},\cdots,\tilde{\bm{G}}^{\lceil\frac{m}{b}\rceil}_{R})\bm{\Psi}_{R}$ and $\bm{P}=\bm{\Psi}_{P}^{\top}\text{Diag}(\tilde{\bm{G}}^{1}_{P},\cdots,\tilde{\bm{G}}^{\lceil\frac{n}{b}\rceil}_{P})\bm{\Psi}_{P}$.
@@ 4 Discussions and Intriguing Insights
AdamW mnmn $3mn$
GaLore [82] mnmn $mn+mr+2nr$
POET (FS) b(b1)b(b-1) mn+3b(b1)mn+3b(b-1)
@@
Performance under a constant parameter budget. Since POET optimizes two orthogonal matrices $\bm{R},\bm{P}$ simultaneously, a natural question arises: which matrix should receive more parameter budget under a fixed total constraint? To investigate this, we conduct a controlled experiment where different ratios of trainable parameters are allocated to $\bm{R}$ and $\bm{P}$ under a fixed total budget. All other settings (e.g., architecture, data) remain unchanged, with full details provided in the Appendix. We use validation perplexity as the evaluation metric. The total parameter budget matches that of fully stochastic POET with $b=\frac{1}{h}m$ for $\bm{R}$ and $b=\frac{1}{h}n$ for $\bm{P}$, where $h=8$, 44, and $3$ correspond to small, medium, and large budgets, respectively. We explore seven allocation settings: $\bm{R}:\bm{P}=1:0$ (i.e., orthogonal training [48, 65, 51]), $0.9:0.1$, $0.75:0.25$, $0.5:0.5$ (i.e., standard POET), $0.25:0.75$, $0.1:0.9$, and $0:1$. Results in Figure 5 show that POET with a balanced allocation between $\bm{R}$ and $\bm{P}$ yields the best performance.
@@
where $\lVert\cdot\rVert$ and $\lVert\cdot\rVert_{F}$ denote spectral and Frobenius norm respectively. Those norms remain invariant when training the network with POET and at initialization they can be bounded with high probability using standard results from random matrix theory (Appendix C). The scale at initialization is typically chosen such that $\bm{W}\in\mathbb{R}^{d\times d}$ satisfies $\|\bm{W}\|=O(1)$ and $\|\bm{W}\|=O(\sqrt{d})$ so that $Q=O_{L}(d)$.
@@
The proof of this lemma can be found in Appendix D. The result extends to Block-stochastic SPO as this is strictly more expressive than fully stochastic SPO. The key idea of the proof is similar to the factorization of orthogonal matrices into a product of Givens rotations. Indeed, by multiplying $\bm{R}^{\top}$ with properly chosen primitive matrices $\bm{G}_{i}$ we can create zeros below the diagonal for one column after another. Note that each $\bm{G}_{i}$ has b(b1)/2b(b-1)/2 parameters while $\bm{R}$ has m(m1)/2m(m-1)/2 parameters, which implies that generally at least $\Omega((m/b)^{2})$ primitives are necessary. In Appendix D we also provide a heuristic that with high probability for $c=O(\ln(m)(m/b)^{2})$ every orthogonal matrix can be written as a product of $c$ orthogonal primitives $\bm{G}_{i}$.
@@
Dynamics of singular spectrum. Inspired by [46], we conduct a spectral analysis by comparing the singular spectrum among AdamW, Muon and POET. Following [1, 68], we compute SVD entropy of the trained Llama-60M model at different iteration. Specifically, the SVD entropy, defined as $H(\bm{\sigma})=\frac{1}{\log n}\sum_{i}\frac{\sigma_{i}^{2}}{\sum_{j}\sigma_{j}^{2}}\log\frac{\sigma_{i}^{2}}{\sum_{j}\sigma_{j}^{2}}$ measures the diversity of singular values; higher entropy indicates a more uniform and diverse spectrum. [46] attributes the favorable performance of Muon over AdamW to the more diverse spectrum of weight matrices updates. As shown in Figure 6, POET consistently maintains high spectral diversity throughout training, owing to its orthogonal equivalence transformation. Therefore, POET can better explore diverse optimization directions.
POET minimizes hyperspherical energy. While POET generalizes energy-preserving training to spectrum-preserving training, it still ensures low hyperspherical energy when initialized with a zero-mean isotropic Gaussian distribution. This is significant, as POET retains the generalization benefits of minimal hyperspherical energy [47, 49, 48]. This property comes from the invariance of zero-mean isotropic Gaussian to orthogonal transformation. [48] shows that for a weight matrix $\bm{W}$ initialized by zero-mean isotropic Gaussian, its neurons, after being normalized, are uniformly distributed on the unit hypersphere. This property provably leads to a small hyperspherical energy. Since zero-mean isotropic Gaussian is invariant to orthogonal transformation, $\text{OET}(\bm{W};\bm{R},\bm{P})=\bm{R}\bm{W}\bm{P}$ does not change the distribution of $\bm{W}$, i.e., $\bm{R}\bm{W}\bm{P}=_{d}\bm{W}$. We give a derivation in Appendix B. Therefore, the transformed neurons of the new weight matrix $\bm{R}\bm{W}\bm{P}$ are also uniformly distributed over the unit hypersphere after being normalized. This validates that POET provably preserves a small energy.
@@
The property that POET preserves singular spectrum while retaining small hyperspherical energy well justifies why POET yields stable training and good generalization. Such a property only holds true when the weight initialization follows zero-mean isotropic Gaussian, which is perfectly satisfied by current weight initialization. It may also justify why zero-mean isotropic Gaussian initialization works better than the uniform spectrum initialization in our experiments. To further validate that POET maintains a small hyperspherical energy, we plot the hyperspherical energy of the Llama-60M model at different iterations in Figure 7. We compare the sum of the hyperspherical energy of all the layers trained by AdamW, Muon or POET. The results again verify POET’s energy-minimizing property. Different from previous orthogonal training [48], POET can not preserve exactly the same hyperspherical energy during training, but it can well minimize hyperspherical energy in an expectation sense. We also find that Muon can better minimize hyperspherical energy than AdamW.
@@ 5.1 LLM Pretraining using LLaMA Transformers
POETBS,bb=64 29.52 (2.39M) 24.52 (5.52M) 20.29 (14.90M) 18.28 (29.22M)
POETBS,bb=128 26.90 (4.81M) 21.86 (11.12M) 18.05 (30.04M) 16.24 (58.91M)
POETBS,bb=256 25.29 (9.66M) 19.88 (22.33M) 16.27 (60.32M) 14.56 (118.26M)
POETFS,bb=1/8 34.06 (0.53M) 29.67 (1.78M) 24.61 (6.34M) 18.46 (25.39M)
POETFS,bb=1/4 28.69 (2.13M) 23.55 (7.13M) 19.42 (25.44M) 17.60 (101.66M)
POETFS,bb=1/2 25.37 (8.54M) 19.94 (28.56M) 15.95 (101.86M) 13.70 (406.88M)
@@
We perform the pretraining experiments on the Llama transformers of varying sizes (60M, 130M, 350M, 1.3B) for POET. We use the C4 dataset [66], a cleaned web crawl corpus from Common Crawl, widely used for LLM pretraining [82, 56, 29]. For POET-BS, $b$ is the block size of the block-diagonal orthogonal matrix. For POET-FS, $b_{\text{in}}$=bmbm for $\bm{R}$ and $b_{\text{out}}$=bnbn for $\bm{P}$. We compare POET against GaLore [82], a low-rank pretraining method, and AdamW, the standard pretraining optimizer. We generally follow the settings in [82]. To better simulate the practical pretraining setting, we significantly increase the number of training tokens for all methods.
Table 2 shows that both POET-FS (bb=1/2) and POET-BS (bb=256) consistently outperform both GaLore and AdamW with significantly fewer parameters. For LLaMA-1B, POET-FS (bb=1/2) yields the best overall performance, achieving a validation perplexity of $13.70$, much better than AdamW ($14.73$) and GaLore ($18.33$). Block-stochastic POET with bb=256 achieves the second-best performance
@@
($14.56$), which still surpasses AdamW with only one-tenth of AdamW’s trainable parameters. Similar patterns can be observed for models of smaller sizes. Moreover, we compare the training dynamics between AdamW and POET in Figure 8. The training dynamics of POET is quite different from AdamW. After an initial rapid drop in perplexity, POET improves more slowly than AdamW. As seen in Phase II (Figure 2), this slower but stable progress can lead to better performance in later stages. We attribute this intriguing phenomenon to the unique reparameterization of POET and How we efficiently approximate orthogonality. The exact mechanism behind this phenomenon remains an open question, and understanding it could offer valuable insights into large-scale model training.
@@
Figure 9: Validation perplexity dynamics of POET (FS, bb=1/2) and AdamW on Llama-60M. POET outperforms the AdamW trained with almost twice the number of seen tokens.
@@
To highlight POET’s non-trivial performance improvement, we increase the training steps (i.e., effectively tokens seen) for AdamW, and find that POET-FS (bb=1/2) still outperforms AdamW even even if AdamW is trained with almost triple the number of tokens. Results are given in Figure 9. In this experiment, the AdamW learning rate was carefully tuned for the full training run, and no training tokens were repeated. Thus, the improvement is non-trivial and cannot be attributed to merely increasing training steps. Interestingly, we also observe from Table 2 that POET’s performance appears strongly correlated with the parameter budget and larger budgets consistently yield better results across model scales. This is particularly important for model scaling law [35]. Another notable observation is that POET significantly outperforms LoRA [26] given a similar parameter budget. For instance, with approximately 30M trainable parameters, POET attains a validation perplexity of $18.05$, significantly better than LoRA’s $25.19$. We further observe that the block-stochastic variant is more parameter-efficient than the fully stochastic one. On the 130M model, it achieves a validation perplexity of $19.88$ with nearly 6M fewer trainable parameters, compared to $19.94$ for the fully stochastic variant. We hypothesize that this is due to better coverage of weight parameters. Specifically, the block-stochastic variant ensures all corresponding weights are updated at each step, unlike the more uneven updates in the fully stochastic variant. Experimental details and results on weight update coverage are provided in Appendix H.
Method Perplexity
AdamW 19.61
POETFS,bb=1/2 16.90
Table 3: Validation perplexity of training a LLaMA 3B model on 5 billion tokens.
Pretraining Llama-3B. To demonstrate the scalability of POET, we apply POET to pretrain Llama with 3B parameters while reusing the same hyperparameters from the 1.3B model. This model was trained with only $1/10$ of the tokens used for the 1.3B model. Compared to Table 2, the slightly higher final perplexity is attributed to the reduced training data (5B tokens). Table 3 shows that POET maintains the performance advantage over AdamW at the 3B scale, consistent with the conclusion drawn from Table 2 for smaller models. Our findings confirm that the benefits of POET are not limited to smaller models but can extend robustly to larger scales.
5.2 LLM Finetuning on Downstream Tasks
FT Model CoLA MNLI MRPC QNLI QQP RTE SST-2 STS-B
Full FT AdamW 0.361 0.658 0.696 0.818 0.829 0.534 0.914 0.880
POET 0.523 0.818 0.824 0.885 0.902 0.661 0.920 0.873
OFT AdamW 0.388 0.774 0.689 0.842 0.867 0.531 0.915 0.762
POET 0.437 0.812 0.740 0.855 0.877 0.538 0.924 0.791
POET AdamW 0.435 0.804 0.806 0.856 0.889 0.653 0.904 0.878
POET 0.505 0.821 0.826 0.892 0.902 0.682 0.931 0.887
Table 4: Downstream performance on GLUE between AdamW- and POET-pretrained models.
To better evaluate models beyond the validation perplexity, we show the results of finetuning the trained model on the GLUE benchmark [75]. This benchmark provides a comprehensive assessment of a model’s language understanding capabilities through a diverse set of downstream tasks. The performance on GLUE serves as an important indicator of the transferability of the model’s learned representations. Performance was measured using accuracy for MNLI, MRPC, QNLI, QQP, RTE, and SST-2; the Matthews Correlation Coefficient for CoLA; and the Pearson Correlation Coefficient for STS-B. For each task, we finetune the models for 10 epochs using a consistent learning rate of 2e-5. The evaluation covers several finetuning strategies: full finetuning (Full FT) with AdamW, orthogonal finetuning (OFT) [65, 51], and finetuning with POET. The results in Table 4 show that the POET-pretrained model consistently outperforms the AdamW-pretrained baseline across all tasks and finetuning methods. The results further validate POET’s generalizability. In this setting, OFT uses significantly less parameters than both Full FT and POET, so OFT generally performs worse. However, the comparison between Full FT and POET is fair, since both methods update the entire model.